OpenStudy (anonymous):

Jack and Jill leave their houses at the same time. Jack is walking to Jill's house and Jill is walking to Jack's house. The distance between their houses is 500 feet. If they each walk at a constant pace and after walking for 5 minutes they meet 20 feet closer to Jack's house than to Jill's house, how much faster is Jill walking? A. They are walking at the same pace. B. 4 feet/minute C. 5 feet/minute D. 8 feet/minute E. It cannot be determined from the information given.

OpenStudy (anonymous):

Are you able to find the speed of Jack and Jill?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok.Speed=Distance Travelled/Time so,for Jack Speed=230/5=46 feet/minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you find the speed of Jill now?

OpenStudy (anonymous):

but how did u get 230?

OpenStudy (anonymous):

ok.Think if they both have same speed,after how much feet from Jack house have they meet?

OpenStudy (anonymous):

250

OpenStudy (anonymous):

250-20= 230

OpenStudy (anonymous):

yes you got it.So what will be Jill's speed?

OpenStudy (anonymous):

54 feet/minute?

OpenStudy (anonymous):

and 54 -46= 8

OpenStudy (anonymous):

great...

OpenStudy (anonymous):

so is it 8 feet/ minute?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Thanks! :D

OpenStudy (anonymous):

@terius7, if you got time I would add something here.

OpenStudy (anonymous):

Yes. What is it?

OpenStudy (anonymous):